1. A TPN order calls for a liter of solution to contain 3.5% of amino acids and 15% of dextrose. How many milliliters each of 8.5% amino acid injection, 70% dextrose injection, and sterile water for injection should be used to prepare the solution?
Answer #1

Amino acids

3.5g \times 1000ml \div 100ml = 35g

35g \times 100ml \div 8.5 g = 411.76ml can be rounded to 412 ml

Dextrose

15g \times 1000ml \div 100ml = 150g

150g \times 100ml \div 70 g = 214 ml

Sterile water for injection

1000ml - 412ml - 214ml = 374 ml